a driver takes 0.20s to apply the breaks after he see a needs for it.this is called reaction time of the driver. if he driving a car at a speed of 54km/hr the breaks cause deacceleration of 6m/s^2 find the distance travalled by the car affter he sees the needs to put the break

Grade:12

1 Answers

Arun
5 years ago
The reaction time of the driver is 0.20 seconds.

so, at the instant the driver feels the need to apply brakes we have to determine the distance covered by the car in that duration.

The speed of car is 54 km/hr = 54 x(5/18) m/s = 15m/s

thus in the period of 0.20 seconds he moves with a constant velocity of 15 m/s before he applies brakes.

so the total distance travelled will be  = v x t

s = v x t

or, s = 15 x 0.20 = 3 m

there after it decelerates to zero m/s

v= 0 m/s

u = 15 m/s

s= ?

using equation of motion:

v2 =u2 +2as

0 = 152 - 2 x 6 x s
or s = 225/12 = 18.75 m

so total distance covered = 3 +18.75 = 21.75 m